9e^2+24e+16=0
It's an important summer assignment & I really need help. Thank you SO much!!

Answer by checkley75(3666)

9e^2+24e+16=0
(3e+4)(3e+4)=0
3e+4=0
3e=-4
e=-4/3 or -1.3333 answer.